OpenStudy (anonymous):

solve the system of equations by graphing. then classify the system x+y=15 x-y=-3

OpenStudy (anonymous):

If you add them together: x+y=15 + x-y=3 ____________ 2x+0=18 2x=18 x = 9

OpenStudy (anonymous):

so would the coordinates be (2,18) and (9,0) ?

OpenStudy (anonymous):

I missed a (-) sign in my addition. It should be 2x+0 = 12.

OpenStudy (anonymous):

Therefore x=6 and y=3

OpenStudy (anonymous):

how do I find the solution for this system of equations?